OpenStudy (anonymous):

find all real solutions sqrt 5x-1 = sqrt 4x+1

OpenStudy (anonymous):

\(f(x) = \sqrt{x}\) is a one to one function. Therefore: \[f(a) = f(b) \implies a = b\] So just equate 5x-1 and 4x+1 and stipulate that for real solutions: \(5x \ge 1\) and \(4x \ge -1\)

OpenStudy (anonymous):

thx